OpenStudy (pokemon23):

Find three consecutive odd integers whose sum is: -27 Find three consecutive odd integers whose sum is: -27 @Mathematics

OpenStudy (anonymous):

x+x+1+x+2=-27

OpenStudy (anonymous):

(x-2) + x + (x +2) = -27 3x = -27 x = -9

OpenStudy (anonymous):

Oh yeah, mine is wrong, didn't notice 'odd' integers

OpenStudy (anonymous):

thus u having three consecutive numbers -7, -9 and -11

OpenStudy (mertsj):

The integers are x x + 2 and x + 4. Their sum is -27 sox+x+2+x+4 = -27 3x+6 = -27 3x = -33 x = -11 x+2=-9 x+4=-7